# The Fragile Nature of Road Transportation Systems

### v2

@@ 3 Problem formulation
We study the onset of disruptions focusing only on the two stable traffic states before and after the disruptions, while the recovery is regarded as a gradual process of congestion dissipation.
For the onset of demand disruptions, we assume a rapid influx of traffic into the network, which can be characterized as an instantaneous event. For the onset of supply disruptions, although it may take some time for the number of vehicles within the network to accumulate, we understand it as a change from one stable traffic state $G(k_{0})$ to the final stable state $(1-r)G(k^{\prime})$, i.e., the equilibrium points in blue as shown in Fig. 4(b). The process of recovery from demand or supply disruptions, on the other hand, takes significantly longer as it is a traffic unloading and congestion dissipation process instead of an instantaneous event.
@@
For demand disruptions, we assume $k^{\prime}>k_{c}$, whereas $k^{\prime}<k_{c}$ for supply disruptions.
A surging demand should be considered a disruption only when its presence leads to a reduction in the network’s maximal possible serviceability, i.e., causing the traffic state to enter the congested zone of the MFD and the flow is below the network capacity. On the other hand, for supply disruption with the constant base demand, if the traffic state can ever surpass the maximal capacity of the disrupted MFD profile, indicating the incoming flow is even greater than the maximal capacity of the network after supply disruptions, then the traffic density will continue to accumulate until the network reaches a full gridlock, and there will be no equilibrium point in this case after a disruption, violating Assumption 1.
@@
For demand disruptions, we assume $q<G(k^{\prime})$, whereas $q<(1-r)G(k_{c})$ for supply disruptions.
Likewise, the necessity of this assumption also lies in the avoidance of gridlock for both demand and supply disruptions. If the base demand is higher than the outgoing flow after a demand disruption or the maximal capacity of the network after a supply disruption, then the traffic state will continue to move to complete gridlock.
@@ 4 Mathematical analysis of the fragility of road transportation systems
In this section, we conduct a mathematical analysis to evaluate the potential fragility of road transportation systems at both microscopic and macroscopic levels. The structure of this section is summarized in Table 1. The reason for not considering the recovery process on the microscopic level is that researchers rarely use FDs to study the recovery from congestion on a link level. And even if we do so, we can also easily get the same conclusion with Proposition 3 and or Proposition 6. In the subsequent study, to investigate the instantaneous disruption onset and between different stable states, the Average Time Spent (ATS) serves as the indicator for the overall performance of a link or a network, since ATS remains constant at any stable traffic state. Conversely, for the examination of disruption recovery, we use Total Time Spent (TTS), as applied in Zhou and Gayah (2021); Chen et al. (2022); Rodrigues and Azevedo (2019), to better reflect the temporal costs for all vehicles in the process, considering that the time spent in the network varies significantly for vehicles entering at different times.
@@
Micro
Macro
Micro
Macro
Onset
Proposition 1
Proposition 2
Proposition 4
Proposition 5
Recovery
Proposition 3
Proposition 6
@@ 4.1 Demand disruption
As outlined in Section 3, the presence of a positive second derivative in performance loss concerning the magnitude of disruption serves as an indication of the transportation system’s fragility, therefore, to illustrate the system’s fragility to demand disruption, we analyze the derivatives of time spent, i.e., ATS for the onset of disruptions or TTS for the recovery process, relative to the initial disruption demand, either represented by disruption density $k^{\prime}$ or an initial disruption demand $n^{\prime}$ when we study the relationship between trip completion and vehicle accumulation. If the system is neither fragile nor antifragile, this approach is expected to yield a linearly growing loss in performance alongside the disruption and zero derivatives. Quoting and in line with the famous statistician George Box, “All models are wrong, but some are useful,” we employ both a numerical and an analytical FD/MFD to investigate the fragile properties under the onset of demand disruption at both microscopic and macroscopic scales.
@@
Road transportation systems are fragile with the onset of demand disruptions on the microscopic level.
@@
We choose the second-degree polynomial FD in Greenshields et al. (1934) as the numerical traffic model and the two-regime linear FD Daganzo (1994) as the analytical traffic model.
For the Greenshields second-degree polynomial FD, the following equations describe traffic in a stable state with $a$ and $b$ being the polynomial coefficients. The traffic flow $q$ and average speed $v$ can be determined as in Eq. 2 and Eq. 3, with the speed-density profile shown in Fig. 5(a).
@@
$G(k)$ $=ak^{2}+bk$ (2)
$v(k)$ $=\frac{q}{k}=ak+b$ (3)
@@
$ATS$ $=\frac{L}{v(k^{\prime})}=\frac{L}{ak^{\prime}+b}$ (4)
$\frac{dATS}{dk^{\prime}}$ $=-aL(ak^{\prime}+b)^{-2}$ (5)
$\frac{d^{2}ATS}{dk^{\prime 2}}$ $=2a^{2}L(ak^{\prime}+b)^{-3}$ (6)
Since the coefficient $a$ is negative and the physical meaning of the term $ak^{\prime}+b$ is the average speed of vehicles on this link, which should always be positive unless in gridlock, thus, both the first and second derivatives of ATS over $k^{\prime}$ are positive, indicating the fragility of transportation systems on a microscopic level.
@@
$v(k)=\begin{cases}v_{f},&0\leq k<k_{c}\\ w+\frac{c}{k},&k_{c}\leq k\leq k_{\rm max}\end{cases}$ (7)
@@
$ATS=\frac{L}{v_{f}}$ (8)
$\frac{dATS}{dk^{\prime}}=0$ (9)
$\frac{d^{2}ATS}{dk^{\prime 2}}=0$ (10)
With the second derivative being zero, it indicates the traffic states before $k_{c}$ are neither fragile nor antifragile. However, as per Assumption 2, the congested area of the MFD is the study focus for demand disruptions, so now we study the derivatives when $k^{\prime}$ is over $k_{c}$:
$ATS$ $=\frac{L}{v(k^{\prime})}=\frac{L}{w+\frac{c}{k^{\prime}}}$ (11)
$\frac{dATS}{dk^{\prime}}$ $=\frac{cL}{(wk^{\prime}+c)^{2}}$ (12)
$\frac{d^{2}ATS}{dk^{\prime 2}}$ $=\frac{-2wcL}{(wk^{\prime}+c)^{3}}$ (13)
Before the disruption density $k^{\prime}$ reaches the maximal density $k_{\rm max}$ of this link, $wk^{\prime}+c>0$ always holds true, and since $w<0$ as well as $c>0$, therefore, both the first and second derivatives are positive.
@@
Road transportation systems are fragile with the onset of demand disruptions on the macroscopic level.
@@
Likewise, on the macroscopic level, we study the fragile property based on the third-degree polynomial MFD in Geroliminis et al. (2013) and multi-regime linear MFD based on MoC in Daganzo and Geroliminis (2008). For MFD approximated with a third-degree polynomial, similar to the Greenshields FD in Eq. 2, we have:
$G(k)$ $=ak^{3}+bk^{2}+ck$ (14)
$v(k)$ $=\frac{q}{k}=ak^{2}+bk+c$ (15)
@@
$ATS$ $=\frac{L}{v(k^{\prime})}=\frac{L}{ak^{\prime 2}+bk^{\prime}+c}$ (16)
$\frac{dATS}{dk^{\prime}}$ $=\frac{-(2ak^{\prime}+b)L}{(ak^{\prime 2}+bk^{\prime}+c)^{2}}$ (17)
$\frac{d^{2}ATS}{dk^{\prime 2}}$ $=\frac{\frac{3}{2}(2ak^{\prime}+b)^{2}+\frac{1}{2}(b^{2}-4ac)}{(ak^{\prime 2}+bk^{\prime}+c)^{3}}L$ (18)
The average speed has to be a real number, indicating Eq. 15 should have real roots, so $b^{2}-4ac>0$ should hold true. Therefore, the derivatives are positive.
The MFD derived with MoC can be approximated by a series of linear functions. Likewise to the Daganzo two-regime linear FD as in Eq. 13, for any linear function, the second derivative is:
@@
As coefficient $c_{i}$ is positive for any cut since the y-intercept should always be positive by definition of MoC, whether the second derivative is positive or negative depends solely on $u_{i}$. The cuts that intercept the MFD before the critical density $k_{c}$ exhibit antifragile properties ($u_{i}>0$, in blue as shown in Fig. 6) while the others with intercepts larger than the critical density $k_{c}$ show fragile responses ($u_{i}<0$, in red), with an exception in case there’s a cut at the critical density ($u_{i}=0$, in gray). Conforming to Assumption 2, for demand disruptions, we focus on the cuts with intercepts larger than the critical density $k_{c}$ ($u_{i}<0$). The second derivative for these cuts is positive indicating a fragile property.
@@
According to Assumption 1, we simplify this surging demand as a disruption that takes place instantly in the network, denoted as $n^{\prime}$ at time $t^{\prime}=0$. As the MoC is composed of a series of linear functions with decreasing gradients as the vehicle accumulation increases, the multi-linear regimes can be represented into multiple sets of consecutive duo linear functions as Fig. 7 shows. The four constants $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ are the slope and y-intercept on the coordinates for the two cuts, with $a_{2}>a_{1}$ and $b_{1}>b_{2}>0$. We refer to these two linear functions as the more congested branch and the less congested branch. Also, the critical accumulation $n_{c}$ here does not represent the critical point of the entire MFD, but rather the critical accumulation of any two consecutive cuts. After a certain period $t_{c}$, the number of vehicles in the network reaches this critical accumulation $n_{c}$. And after any period $t>t_{c}$, the vehicle accumulation becomes $n$. We also denote the initial trip completion and critical trip completion as $m_{0}=a_{1}n^{\prime}+b_{1}$ and $m_{c}=a_{1}n_{c}+b_{1}=a_{2}n_{c}+b_{2}$ respectively.
@@
Any two consecutive cuts of the MFD can be formulated into the following Eq. 20:
$G(n)=\begin{cases}a_{1}n+b_{1},&n_{c}\leq n<n_{\rm max}\\ a_{2}n+b_{2},&0\leq n<n_{c}\end{cases}$ (20)
The system dynamics can be summarized as the following Eq. 21.
$\frac{dn}{dt}=-G(n)+q=-a_{i}n-b_{i}+q$ (21)
When the traffic states move only along a single branch, and given any initial vehicle accumulation $n_{1}$ at the beginning of a period from $t_{1}$ to $t_{2}$, the number of vehicles $n_{2}$ at the end of this period can be determined as:
@@
$t_{2}-t_{1}$ $=-\frac{1}{a_{i}}\ln{\left(\frac{a_{i}n_{2}+b_{i}-q}{a_{i}n_{1}+b_{i}-q}\right)}$ (23)
$n_{2}$ $=\frac{e^{-a_{i}(t_{2}-t_{1})}(a_{i}n_{1}+b_{i}-q)}{a_{i}}-\frac{b_{i}-q}{a_{i}}$ (24)
Therefore, with the disruption accumulation $n^{\prime}$, and when the traffic states are on the same branch. After any time $t$, the vehicle accumulation $n$ would be:
$n$ $=\frac{a_{1}n^{\prime}+b_{1}-q}{a_{1}}e^{-a_{1}t}-\frac{b_{1}-q}{a_{1}}$ (25)
The TTS in this case can be calculated as:
@@
$=-\frac{a_{1}n^{\prime}+b_{1}-q}{a_{1}^{2}}e^{-a_{1}t}-\frac{b_{1}-q}{a_{1}}t+\frac{a_{1}n^{\prime}+b_{1}-q}{a_{1}^{2}}$ (27)
Now we calculate the derivatives of TTS considering $t$ as any positive constant.
$\frac{dTTS}{dn^{\prime}}$ $=\frac{1}{a_{1}}-\frac{e^{-a_{1}t}}{a_{1}}$ (28)
$\frac{d^{2}TTS}{dn^{\prime 2}}$ $=0$ (29)
The second derivative of TTS is 00, indicating that when the traffic states move only along a single branch, it shows neither fragility nor antifragility.
On the other hand, when the traffic state goes over the critical vehicle accumulation $n_{c}$, and since the MoC is a piecewise function, we calculate the TTS separately on both the more congested and the less congested branches, denoted as $TTS_{1}$ and $TTS_{2}$. Since the critical time $t_{c}$ is still unknown, we need to determine $t_{c}$ first, similar to Eq. 23.
$t_{c}$ $=-\frac{1}{a_{1}}\ln{\left(\frac{a_{1}n_{c}+b_{1}-q}{a_{1}n^{\prime}+b_{1}-q}\right)}$ (30)
As both $a_{1}n_{c}+b_{1}$ and $a_{2}n_{c}+b_{2}$ are equal to $m_{c}$, we can rewrite the above Eq. 30 as:
$t_{c}$ $=-\frac{1}{a_{1}}\ln{\left(\frac{m_{c}-q}{a_{1}n^{\prime}+b_{1}-q}\right)}$ (31)
Likewise to Eq. 27, the $TTS_{1}$ for the more congested branch is:
$TTS_{1}$ $=-\frac{a_{1}n^{\prime}+b_{1}-q}{a_{1}^{2}}e^{-a_{1}t_{c}}-\frac{b_{1}-q}{a_{1}}t_{c}+\frac{a_{1}n^{\prime}+b_{1}-q}{a_{1}^{2}}$ (32)
$=-\frac{m_{c}-q}{a_{1}^{2}}+\frac{b_{1}-q}{a_{1}^{2}}\ln{\left(\frac{m_{c}-q}{a_{1}n^{\prime}+b_{1}-q}\right)}+\frac{a_{1}n^{\prime}+b_{1}-q}{a_{1}^{2}}$ (33)
Since TTS is the sum of $TTS_{1}$ and $TTS_{2}$, the second derivative of TTS would also be the sum of the derivatives. The derivatives for $TTS_{1}$ are:
$\frac{dTTS_{1}}{dn^{\prime}}$ $=-\frac{b_{1}-q}{a_{1}}(a_{1}n^{\prime}+b_{1}-q)^{-1}+\frac{1}{a_{1}}$ (34)
$\frac{d^{2}TTS_{1}}{dn^{\prime 2}}$ $=(b_{1}-q)(a_{1}n^{\prime}+b_{1}-q)^{-2}$ (35)
According to Eq. 24, the vehicle accumulation on the less congested branch from $t_{c}$ to $t$ would be:
$n$ $=\frac{e^{-a_{2}(t-t_{c})}(a_{2}n_{c}+b_{2}-q)}{a_{2}}-\frac{b_{2}-q}{a_{2}}$ (36)
$=\frac{e^{a_{2}t_{c}}(m_{c}-q)}{a_{2}}e^{-a_{2}t}-\frac{b_{2}-q}{a_{2}}$ (37)
The $TTS_{2}$ for the less congested branch would be:

### v4

@@ 3 Problem formulation
We study the onset of disruptions focusing only on the stable traffic state after the disruptions, while the recovery is regarded as a gradual process of congestion dissipation.
For the onset of demand disruptions, we assume a rapid influx of traffic into the network, which can be characterized as an instantaneous shift of the equilibrium point in blue as shown in Fig. 4(a). For the onset of supply disruptions, although it may take some time for the number of vehicles within the network to accumulate, we also understand it as an instantaneous change from the original stable traffic state $q_{0}=G(k_{0})$ to the new stable state $q_{0}=(1-r)G(k^{\prime})$ after disruption. On the other hand, the recovery process from demand or supply disruptions takes significantly longer as it is a traffic unloading and congestion dissipation process instead of an instantaneous event.
@@
We assume the disruption density is greater than the critical density $k^{\prime}>k_{c}$ for demand disruptions, whereas $k^{\prime}<k_{c}$ for supply disruptions.
A surging demand should be considered a disruption only when the traffic state $(k^{\prime},G(k^{\prime}))$ enters the congested zone of the MFD with $k^{\prime}>k_{c}$, causing a reduction in the network’s maximal possible serviceability. On the other hand, for supply disruption under the constant base demand $q_{0}$, if the traffic state can ever surpass the maximal capacity of the disrupted MFD profile, i.e., $k^{\prime}>k_{c}$, indicating that $q_{0}$ has to be greater than the disrupted maximal capacity of the network $(1-r)q_{\rm max}$. In this case, the traffic density will continue to accumulate until the network reaches a full gridlock, and there will be no equilibrium point in this case after a disruption, violating Assumption 1.
@@
We assume the base demand $q_{0}<G(k^{\prime})$ for demand disruption, whereas $q_{0}<(1-r)G(k_{c})$ for supply disruptions.
Likewise, this assumption also lies in avoiding gridlock for both demand and supply disruptions. If the base demand $q_{0}$ is higher than the outflow $G(k^{\prime})$ after a demand disruption or the maximal capacity of the network $(1-r)G(k_{c})$ after a supply disruption, then the traffic state will continue to move to complete gridlock $(k_{\rm max},0)$.
@@ 4 Mathematical analysis of the fragility of road transportation systems
In this section, we conduct a mathematical analysis to evaluate the potential fragility of road transportation systems at both link and network levels. The structure of this section is summarized in Table 2. The reason for not considering the recovery process at the link level is that researchers rarely use FDs to study the recovery from congestion for a link section. And even if we do so, we can also easily get the same conclusion with Proposition 3 and Proposition 6. In the subsequent study, to investigate the instantaneous disruption onset between different stable states, the Average Time Spent (ATS) serves as the indicator for the overall performance of a link or a network, since ATS remains constant at any stable traffic state. Conversely, for the examination of disruption recovery, we use Total Time Spent (TTS), as applied in Rodrigues and Azevedo (2019); Zhou and Gayah (2021); Chen et al. (2022), to better reflect the temporal costs for all vehicles in the process, considering that the time spent in the network varies significantly for vehicles entering at different times.
@@
Link Network Link Network
Onset Proposition 1 Proposition 2 Proposition 4 Proposition 5
Recovery - Proposition 3 - Proposition 6
@@ 4.1 Demand disruption
As outlined in Section 3, the presence of a positive second derivative in performance loss regarding the magnitude of disruption serves as an indication of the transportation system’s fragility. Therefore, to illustrate the system’s fragility to demand disruption, we analyze the derivatives of time spent, namely ATS for the onset of disruptions or TTS for the recovery process, relative to the initial disruption demand, either represented by disruption density $k^{\prime}$ or an initial disruption demand $n^{\prime}$ when we study the relationship between trip completion and vehicle accumulation. If the system is neither fragile nor antifragile, this approach is expected to yield a linearly growing loss in performance alongside the disruption and zero derivatives. Regarding the traffic models to be selected, numerical forms of FD/MFD have the strength of generally being considered more accurate as they are directly approximated from mass sensor data with algebraic functions such as polynomials or exponential functions. On the other hand, analytical forms of FDs/MFD are derived from parameters with physical meanings like capacity and free flow speed without the presence of sensor data, allowing for the estimate of the traffic model with the least amount of information and costs for sensor installation. Therefore, we employ both a numerical and an analytical FD/MFD to investigate the fragile properties under the onset of demand disruption at the link level and the whole network.
@@
Road transportation systems are fragile with the onset of demand disruptions at the link level.
@@
We choose the second-degree polynomial FD in Greenshields et al. (1934) as the numerical traffic model and the two-regime linear FD in Daganzo (1994) as the analytical traffic model.
For the Greenshields second-degree polynomial FD, the following equations describe traffic in a stable state with $\alpha_{1}$ and $\alpha_{2}$ being the polynomial coefficients. The traffic flow $q$ and average speed $v$ can be determined as in Eq. 3a and Eq. 3b, with the speed-density profile shown in Fig. 5(a).
@@
$G(k)$ $=\alpha_{2}k^{2}+\alpha_{1}k$ (3a)
$v(k)$ $=\frac{q}{k}=\alpha_{2}k+\alpha_{1}$ (3b)
@@
$ATS$ $=\frac{l}{v(k^{\prime})}=\frac{l}{\alpha_{2}k^{\prime}+\alpha_{1}}$ (4a)
$\frac{dATS}{dk^{\prime}}$ $=-\alpha_{2}(\alpha_{2}k^{\prime}+\alpha_{1})^{-2}l$ (4b)
$\frac{d^{2}ATS}{dk^{\prime 2}}$ $=2\alpha_{2}^{2}(\alpha_{2}k^{\prime}+\alpha_{1})^{-3}l$ (4c)
Since the coefficient $\alpha_{2}$ is negative and the physical meaning of the term $\alpha_{2}k^{\prime}+\alpha_{1}$ is the average speed of vehicles at this link, which should always be positive until reaching gridlock, both the first and second derivatives of ATS over $k^{\prime}$ are positive, indicating the fragility of transportation systems at a link level.
@@
$v(k)=\begin{cases}u_{f},&0\leq k<k_{c}\\ w+\frac{c_{1}}{k},&k_{c}\leq k<k_{\rm max}\end{cases}$ (5)
@@
$ATS=\frac{l}{u_{f}}$ (6a)
$\frac{dATS}{dk^{\prime}}=0$ (6b)
$\frac{d^{2}ATS}{dk^{\prime 2}}=0$ (6c)
The second derivative being zero indicates that the traffic states before $k_{c}$ are neither fragile nor antifragile. However, as per Assumption 2, the congested area of the MFD is the study focus for demand disruptions, so now we study the derivatives when $k^{\prime}$ is over $k_{c}$:
$ATS$ $=\frac{l}{v(k^{\prime})}=\frac{l}{w+\frac{c_{1}}{k^{\prime}}}$ (7a)
$\frac{dATS}{dk^{\prime}}$ $=\frac{c_{1}l}{(wk^{\prime}+c_{1})^{2}}$ (7b)
$\frac{d^{2}ATS}{dk^{\prime 2}}$ $=\frac{-2wc_{1}l}{(wk^{\prime}+c_{1})^{3}}$ (7c)
Before the disruption density $k^{\prime}$ reaches the maximal density $k_{\rm max}$ of this link, $wk^{\prime}+c_{1}>0$ always holds true, and since $w<0$ as well as $c_{1}>0$, therefore, both the first and second derivatives are positive.
@@
Road transportation systems are fragile with the onset of demand disruptions at the network level.
@@
Likewise, at the network level, we study the fragile property based on the third-degree polynomial MFD in Geroliminis et al. (2013) and multi-regime linear MFD based on MoC in Daganzo and Geroliminis (2008). For MFD approximated with a third-degree polynomial, similar to the Greenshields FD in Eq. 3a, we have:
$G(k)$ $=\alpha_{3}k^{3}+\alpha_{2}k^{2}+\alpha_{1}k$ (8a)
$v(k)$ $=\frac{q}{k}=\alpha_{3}k^{2}+\alpha_{2}k+\alpha_{1}$ (8b)
@@
$ATS$ $=\frac{l}{v(k^{\prime})}=\frac{l}{\alpha_{3}k^{\prime 2}+\alpha_{2}k^{\prime}+\alpha_{1}}$ (9a)
$\frac{dATS}{dk^{\prime}}$ $=\frac{-(2\alpha_{3}k^{\prime}+\alpha_{2})l}{(\alpha_{3}k^{\prime 2}+\alpha_{2}k^{\prime}+\alpha_{1})^{2}}$ (9b)
$\frac{d^{2}ATS}{dk^{\prime 2}}$ $=\frac{\frac{3}{2}(2\alpha_{3}k^{\prime}+\alpha_{2})^{2}+\frac{1}{2}(\alpha_{2}^{2}-4\alpha_{1}\alpha_{3})}{(\alpha_{3}k^{\prime 2}+\alpha_{2}k^{\prime}+\alpha_{1})^{3}}l$ (9c)
The average speed has to be a real number, indicating Eq. 8b should have real roots, so $\alpha_{2}^{2}-4\alpha_{1}\alpha_{3}>0$ should hold true. Thus, since $\alpha_{3}k^{\prime 2}+\alpha_{2}k^{\prime}+\alpha_{1}$ represents the speed and is therefore positive, the derivatives are also positive.
The MFD derived from MoC can be approximated with a series of linear functions. Likewise to the Daganzo two-regime linear FD as in Eq. 7c, for any linear function, the second derivative is:
@@
As coefficient $c_{i}$ is non-negative for any cut $i$ since the y-intercept should always be non-negative by definition of MoC, whether the second derivative is positive or negative depends solely on $u_{i}$. The cuts that intercept the MFD before the critical density $k_{c}$ exhibit antifragile properties ($u_{i}>0$, in blue as shown in Fig. 6) while the others with intercepts larger than the critical density $k_{c}$ show fragile responses ($u_{i}<0$, in red), with an exception in case there’s a cut at the critical density ($u_{i}=0$, in gray). Conforming to Assumption 2, for demand disruptions, we focus on the cuts with intercepts larger than the critical density $k_{c}$ ($u_{i}<0$). The second derivative for these cuts is positive indicating a fragile property.
@@
The MoC is composed of a series of linear functions, which are referred to as cuts $1,2,\dots,N$, with increasing gradients as the vehicle accumulation decreases, and such multi-linear regimes are presented in Fig. 7, where the black dashed lines indicate multiple cuts for simplicity. For a given cut $i$, the slope and the y-intercept on the coordinates are denoted as $a_{i}$ and $b_{i}$, with $a_{i+1}>a_{i}$ and $b_{i}>b_{i+1}>0$. The critical accumulation $n_{\tilde{c},i}$ on the $x$-axis does not represent the critical point of the entire MFD, but rather the critical accumulation between any two consecutive cuts $i$ and $i+1$ with two exceptional cases being $n_{\tilde{c},0}=n_{\rm max}$ and $n_{\tilde{c},N}=0$. According to Assumption 1, we simplify this surging demand as a disruption that takes place instantly in the network, denoted as $n^{\prime}$ at time $t_{0}=0$ and lands on cut $y$. At time $t_{\tilde{c},i}\;(y\leq i<z)$, the number of vehicles in the network reaches this critical accumulation $n_{\tilde{c},i}$. And after any period $t$, the traffic state lands on the cut $z$. A virtual intersection point $(n_{\check{c}},m_{\check{c}})$ between cut $y$ and cut $z$ is also illustrated and will be discussed later in this section. We also denote the initial trip completion and critical trip completion as $m^{\prime}=a_{y}n^{\prime}+b_{y}$ and $m_{\tilde{c},i}=a_{i}n_{\tilde{c},i}+b_{i}=a_{i+1}n_{\tilde{c},i}+b_{i+1}$ respectively.
@@
Any cut $i$ of the MFD can be formulated into the following Eq. 11:
$M(n)=a_{i}n+b_{i},\forall n\in(n_{\tilde{c},i},n_{\tilde{c},i-1}]$ (11)
By calculating the difference between the base demand $m_{0}$ as the inflow rate and the trip completion $M(n)$ as the outflow rate, the system dynamics on cut $i$ can be summarized as Eq. 12:
$\frac{dn}{d\tau}=-M(n)+m_{0}=-a_{i}n-b_{i}+m_{0}$ (12)
Assuming that the traffic states move only along a single cut $i$, and with any amount of vehicle accumulation $n_{1}$ at the beginning of a given period between $t_{1}$ and $t_{2}$, the number of vehicles $n_{2}$ at the end of this period $t$ can be determined as: